OpenStudy (anonymous):

divide -2x^3 -4x^2 + 3x +2 by x=-3

OpenStudy (anonymous):

Do you mean substitute x=-3 into the expression? You cannot divide by x=-3

OpenStudy (anonymous):

yes substitute

OpenStudy (anonymous):

So you just need to replace all of the xes with -3: = -2x^3 -4x^2 + 3x +2 = -2*(-3^3) - 4*(-3^2) + 3*-3 + 2 = -2*-27 - 4*9 - 9 + 2 (do the indices) = 54 - 36 - 9 + 2 (last multiplication) = 11

OpenStudy (anonymous):